| <p>06:00</p>
<p>FEE modules power cycled.</p>
<p>No issues on power cycle restart.</p>
<p>All system check OK - clock status, ADC calibration etc.</p>
<p>ASIC settings loaded are the same as previous startup. Slow comparator thresholds 0x32 (500 keV) and all FEEs set to LEC.</p>
<p>Statistics (ADC data) - attachment 1. Note aida10 fluctuations between 10,000 and 300 which appears to be a physics effect (beam on target)</p>
<p>Spectrum browser - layout 1 - attachment 2</p>
<p>Merger - attachment 3</p>
<p>Tape server - attachment 4 Data rate is about 300 kB/sec.</p>
<p>FEE temperatures OK</p>
<p>Waveforms - attachments 4,5</p>
<p>All Lekagae currents appear normal</p>
<p>Online monitoring scropt started. Beam on target. Run 22 started. See nect elog for details.</p> |